Suppose P and Q' are subsets of U. n(U) = 15, n(P) = 6, and n(Q') = 4. Find
a. n(p')
b. n(Q)
I'm quite lost on how to find these. I do understand what we're given, but don't know how to apply it.

Answer by ikleyn(52803)

.
Suppose P and Q' are subsets of U. n(U) = 15, n(P) = 6, and n(Q') = 4. Find
a. n(p')
b. n(Q)
I'm quite lost on how to find these. I do understand what we're given, but don't know how to apply it.
~~~~~~~~~~~~~~

(a)  n(p') = n(U) - n(P) = 15 - 6 = 9.


(b)  n(Q) = n(U) - n(Q') = 15 - 4 = 11.

Solved and answered.

I do not go into further explanations, since you say that you do understand the subject.